# Leveraging Fine-Grained Information and Noise Decoupling for Remote Sensing Change Detection

### v1

@@ 1. Introduction
Figure 1. Conceptual illustration of proposed approach. (a) Two types of change detection pseudo-change. (b) Multi-scale fused features (far left) visualised with features learned by FINO (middle), with FINO predictions on the far right. (c) The core process of FINO. Context-dependent learning (DCL) compensates for fine-grained features, brightness-aware and shape-aware (BAS) perception to decouple task-agnostic noise, and regularization gate (REGA) decouples task-specific noise.
@@
  • 4. We achieve state-of-the-art performance on multiple change detection benchmarks.
@@ 3.1. Overall Framework
In this section we explain our proposed method which addresses the problem of loss of fine-grained information during network downsampling and the difficulty of removing irrelevant change noise in the change detection problem. The framwork of our methods is shown in Fig.  2.
@@
Figure 3. FINO Detail Structure.
@@ 3.2. Context-Dependent Learning
(3) $A_{i}(C_{i},Z_{i})$ $=Softmax(\frac{flatten(\varphi_{1}(C_{i},Z_{i}))}{\sqrt{dk}})$
@@
where $C_{i}$ denotes the difference feature of the Stage-$i-th$ layer in Fig. 3, $T_{i}$ denotes the region-aware contextual features learned by CDL and $\varphi_{1},\varphi_{2}$ denotes the convolution function.
Since change is often of a localised nature, excessive focus on global attention is not effective. SAM is a local attention that differs from the normal attention. Self-attention uses a attention vector multiplication, while we use element multiplication to compute the attention map.
@@ 3.3. Brightness-Aware and Shape-Aware Learning
Shape-aware learning. To refine boundary and expand the receptive field, we use a shape-aware module that consists of mutiple asymmetric convolution for anti-aliasing of the edge, shown as in Fig. 3. In our approach, we use seven convolution kernels, including $1\times 1,1\times 3,3\times 1,3\times 3,1\times 5,5\times 1,5\times 5$. We use Eq to represent the learning of the shape-aware module:
(8) $H_{i}$ $=Asym(T_{i})$
@@
where $Asym$ indicates the asymmetric convolutions and $MLP$ represents a multilayer perceptron mapping.
@@ 3.4. Regularization Gate Structure
(11) $G_{i}$ $=\sigma(\varphi(CAT(X_{t1,i},X_{t2,i},H_{i})))+M_{i}I_{t,i}$ $=X_{t,i}\otimes G_{i}$
@@
We use the equation 12 and 13 to refer the above operations.
@@
where $w$ denotes the weight that maps the features of bitemproal images to the same distribution space. We then obtain a pixel-by-pixel distribution of distances and input them to the segmentation head to get the prediction.
@@ 3.5. Loss
where $\lambda$ indicates a loss weight hyperparameter. We adopt binary cross-entropy for $\mathcal{L}_{cd}$ and $\mathcal{L}_{sal}$.
@@ 4.2. Comparison With State-of-the-Art Methods
The specific results are shown in the Table 3. The Tabel 3 visualizes the considerable improvement of our method over methods such as feature fusion. On both F1 and IoU performing better than DARNet (Li et al., 2022b) which is the the SOTA, where F1 is 0.43 higher than DARNet and IoU is 0.83 higher than DARNet.
@@
4.3. Comparison with other Feature Fusion and Enhancement methods
To compare the superiority of FINO for feature learning at different scales, we compare it with several methods of recent two years that perform well on the LEVIR-CD dataset, which apply feature concatenation (Li et al., 2022a), feature interaction (Lei et al., 2022; Li et al., 2022b), skip connection (Wan et al., 2022), and attention mechanisms (Bandara and Patel, 2022) to achieve feature fusion of different scales.
The specific results are shown in the Table 3. The Tabel 3 visualizes the considerable improvement of our method over methods such as feature fusion. On both F1 and IoU performing better than DARNet (Li et al., 2022b) which is the the SOTA, where F1 is 0.43 higher than DARNet and IoU is 0.83 higher than DARNet.
Methods Pre.(%) Rec.(%) F1(%) IoU(%)
RSCD 92.15 89.73 90.93 83.36
D-TNet18 89.13 88.58 88.85 79.94
DARNet 92.67 91.31 91.98 85.16
ChangeFormer 92.05 88.80 90.40 82.48
DESSN 90.99 91.73 91.36 -
FINO(Ours) 93.70 91.15 92.41 85.89
@@ 4.4. Detection Result
As shown in Figure 6, our method also outperforms other methods in complex scenes. Our method is more robust to noise and provides more accurate detection results. The buildings are highly similar to their surroundings, as shown in the first and second rows of the Figure 6. And the buildings are very similar in shape and colour to the surrounding bush noise, especially as shown in the the red boxes in the second rows and the third rows. Comparing columns (c), (h), and (i), it is clear that our method performs better. Comparing columns (d)-(g) and (i), there are few targets missed. However they detect some noise as a change (e.g. the blue box) because they are oversensitive about the change. Since feature elimination and aggregation module (FEAM) is designed to eliminate noise and robustly enhance the foreground, our method detects all the changed targets without being disturbed by these noises. This demonstrates the high robustness of our approach.
Replace Pre.(%) Rec.(%) F1(%) IoU(%)
Concatenation 92.53 89.79 91.14 83.72
Cross-attention 88.85 81.37 84.94 73.82
No-local attention 90.80 88.64 89.75 81.41
SPP attention 91.97 89.44 90.69 82.96
CDL 93.70 91.15 92.41 85.89
@@ 4.5. Ablation Study and Further Analysis
4.5.1. Importance of the Main Components
To verify the usefulness of our FINO for information recovery, we conducted ablation experiments for the positions of the cascade structure, where the network layers are numbered 1, 2, 3 and 4 from the lower to the higher levels in Fig. 2. The specific results are shown in Table 5. As can be seen in Table 5, the accuracy of change detection are greatly improved as the more layers employed FINO. In particular, during the process from Stage-4 to Stage-2, the improvement of FINO is significant.
@@ 4.5.1. Importance of the Main Components
4.5.2. Importance of Fine-Grained Information
To verify the effectiveness of CDL and proof of significance of fine-grained information, we replace CDL with several other types of attention, including feature concatenation, cross-attention, no-local attention and spatial pyramid pooling (SPP) attention. And due to space constraints, specific structural details are provided in the supplementary material.

### v3

@@ 1. Introduction
Figure 1. Conceptual illustration of proposed approach. (a) Two types of change detection pseudo-change. (b) Multi-scale fused features (far left) visualised with features learned by FINO (middle), with FINO predictions on the far right. (c) The core process of FINO. Context-dependent learning (CDL) compensates for fine-grained features, brightness-aware and shape-aware (BSA) perception to decouple task-agnostic noise, and regularization gate (REGA) decouples task-specific noise.
@@
  • 4. We achieve SOTA performance on multiple change detection benchmarks. Compared to the SOTA methods, our method improves F1 by 0.60% and IoU by 1.03% on the LEVIR-CD (Chen and Shi, 2020) dataset. On the WHU (Ji et al., 2019) dataset, the F1 is improved by 0.86% and IoU by 4.97%. On the DSIFN-CD (Zhang et al., 2020a) dataset, our approach leads to a 1.48% improvement in F1 and a 2.39% increase in IoU. Finally, on the CDD (Lebedev et al., 2018) dataset, we saw an increase of 1.17% in F1 and 2.25% in IoU.
@@ 3.1. Overall Framework
In this section we explain our proposed method which addresses the problem of loss of fine-grained information during network downsampling and the difficulty of removing irrelevant change noise in the change detection problem. The framework of our methods is shown in Fig.  2.
@@
Figure 3. FINO Detail Structure. FINO consists of CDL, shape-aware module and REGA. First, CDL compensates the fine-grained information in $C_{i}$ through $Z_{i+1}$ to obtain the compensated feature $T_{i}$. Then, the shape-aware module learns the shape to obtain $M_{i}$ and enriched feature $H_{i}$. In REGA, $H_{i}$ and the bitemporal features $X_{t_{1},i}X_{t_{2},i}$ pass through a gated structure to obtain the change feature $Z_{i}$ of the current layer, where $M_{i}$ is used as a regularization term for the gated structure to improve the robustness to pseudo changes.
@@ 3.2. Context-Dependent Learning
(3) $A_{i}(C_{i},Z_{i+1})$ $=Softmax(\frac{flatten(\varphi_{1}(C_{i},Z_{i+1}))}{\sqrt{d_{k}}})$
@@
where $d_{k}$ represents the feature dimension, $C_{i}$ denotes the difference feature of the Stage-$i-th$ layer in Fig. 3, $T_{i}$ denotes the region-aware contextual features learned by CDL and $\varphi_{1},\varphi_{2}$ denotes the convolution function.
Since change is often of a localised nature, excessive focus on global attention is not effective. CDL is a local attention that differs from the normal attention. Self-attention uses a attention vector multiplication, while we use element multiplication to compute the attention map.
@@ 3.3. Brightness-Aware and Shape-Aware Learning
Shape-aware learning. To refine boundary and expand the receptive field, we use a shape-aware module that consists of multiple asymmetric convolution for anti-aliasing of the edge, shown as in Fig. 3. In our approach, we use seven convolution kernels, including $1\times 1,1\times 3,3\times 1,3\times 3,1\times 5,5\times 1,5\times 5$. We use Eq. 8 to represent the learning of the shape-aware module:
(8) $H_{i}$ $=T_{i}+W_{0}Asym(T_{i})$
@@
where $Asym$ indicates the asymmetric convolutions, $W_{0}$ denotes the weight of the $1\times 1$ convolution and $MLP$ represents a multilayer perceptron mapping.
@@ 3.4. Regularization Gate Structure
(11) $G_{i}=\sigma(\varphi(CAT(X_{t1,i},X_{t2,i},H_{i})))+M_{i}$
(12) $I_{t,i}=X_{t,i}\otimes G_{i}$
@@
We use the equation 13 and 14 to refer the above operations.
@@
where $w$ denotes the weight of a $1\times 1$ convolution that maps the features of bitemproal images to the same distribution space. We then obtain a pixel-by-pixel distribution of distances and input them to the segmentation head to get the prediction.
@@ 3.5. Loss
where $\lambda$ indicates a loss weight hyperparameter. We adopt binary cross-entropy for $\mathcal{L}_{cd}$ and $\mathcal{L}_{sal}$. Based on task relevance, the change detection loss $\mathcal{L}_{cd}$ and shape-aware learning loss $\mathcal{L}_{sal}$ are the primary losses. The global contrast learning loss $\mathcal{L}_{gcl}$ and region-level alignment loss $\mathcal{L}_{rcl}$ are auxiliary losses, with the weight $\lambda$ set to 0.1.
@@ 4.2. Comparison With State-of-the-Art Methods
The specific results are shown in the Table 2. The Table 2 visualizes the considerable improvement of our method over methods such as feature fusion. On both F1 and IoU performing better than DARNet (Li et al., 2022b) which is the the SOTA, where F1 is 0.43 higher than DARNet and IoU is 0.83 higher than DARNet.
@@ 4.3. Detection Result
As shown in Figure 5, our method also outperforms other methods in complex scenes. Our method is more robust to noise and provides more accurate detection results. The buildings are highly similar to their surroundings, as shown in the first and second rows of the Figure 5. And the buildings are very similar in shape and colour to the surrounding bush noise, especially as shown in the the red boxes in the second rows and the third rows. Comparing columns (c), (h), and (i), it is clear that our method performs better. Comparing columns (d)-(g) and (i), there are few targets missed. However they detect some noise as a change (e.g. the blue box) because they are oversensitive about the change. Since REGA is designed to eliminate noise and robustly enhance the foreground, our method detects all the changed targets without being disturbed by these noises. This demonstrates the high robustness of our approach.
@@ 4.4. Ablation Study and Further Analysis
Importance of the Main Components To verify the usefulness of our FINO for information recovery, we conducted ablation experiments for the positions of the cascade structure, where the network layers are numbered 1, 2, 3 and 4 from the lower to the higher levels in Fig. 2. The specific results are shown in Table 3. As can be seen in Table 3, the accuracy of change detection are greatly improved as the more layers employed FINO. In particular, during the process from Stage-4 to Stage-2, the improvement of FINO is significant.
@@
Importance of Fine-Grained Information To verify the effectiveness of CDL and proof of significance of fine-grained information, we replace CDL with several other types of attention, including feature concatenation, cross-attention, no-local attention and spatial pyramid pooling (SPP) attention. And due to space constraints, specific structural details are provided in the supplementary material.
@@
Replace Pre.(%) Rec.(%) F1(%) IoU(%)
Concatenation 92.53 89.79 91.14 83.72
Cross-attention 88.85 81.37 84.94 73.82
No-local attention 90.80 88.64 89.75 81.41
SPP attention 91.97 89.44 90.69 82.96
CDL 93.70 91.15 92.41 85.89